The volume V of a right circular cone varies jointly as the square of the radius R and the height H. Determine what happens to V when R is six times as large.
Help!
thanks
cz

Answer by stanbon(75887)

The volume V of a right circular cone varies jointly as the square of the radius R and the height H.
---
V = k(R^2*H)
---------------------
Determine what happens to V when R is six times as large.
Let r = 6R
V = k(r^2*H)
V = k((6R)^2*H)
V = k(36R^2*H)
V = 36k(R^2*H)

--------------------
Ans: The volume is multiplied by 36
=======================================
Cheers,
Stan H.
==================